OpenStudy (juan1857):

Find (1-i)^6. Express the result in rectangular form.

OpenStudy (john_es):

Do you know how to expand the power?

OpenStudy (john_es):

You can also change to polar form, then do the sixth power, and came back to the rectangular form.

OpenStudy (juan1857):

I don't know neither ways, can you explain the most simple way?

OpenStudy (john_es):

The most simple way is the polar form, \[z=r_\theta \] As the power is, \[z^n=r^n_{n\theta}\]

OpenStudy (john_es):

And it is related to rectangular form, \[z=r(\cos\theta+i\sin\theta)\]

OpenStudy (john_es):

First find r \[r=|z|=\sqrt{1^2+1^2}=\sqrt{2}\] Then find the argument \[\theta=\arctan(Im(z)/Re(z))=\arctan(-1)=225º\]

OpenStudy (john_es):

Now, we have, \[z=\sqrt{2}_{225}\]So the power is, \[z^6=(\sqrt{2})^6_{6\cdot225}=8_{1350}\]Now find the rectangular form, \[z=8(\cos(1350)+i\sin(1350 ))=-8i\]

OpenStudy (john_es):

That should be the answer if my lines are correct.

OpenStudy (juan1857):

Oh ok thanks again

OpenStudy (john_es):

Sorry, it should be \[z=8i\] because the correct angle was 315 not 225.

OpenStudy (juan1857):

ok thanks!